Aricuzyas:

The price of a notebook was 3.30 yesterday. Today, the price rose to 3.80 . Find the percentage increase. Round your answer to the nearest tenth of a percent. @Convert I need your help

Convert:

Well the formula to find the percent increase is to divide the increase by the original number then multiply by 100: 3.80-3.50= 0.30 0.30 x 100 = 30% increase

Aricuzyas:

Thanks :D

Convert:

No problem :)

Aricuzyas:

I'm suppose to round to the nearest tenths

Convert:

Oh wait I did it wrong-

Convert:

3.80-3.30 = 0.50 0.50/3.30 = 0.1515 0.1515 * 100 = 15.15% increase

Convert:

15.15% rounds up to 15.2%

Aricuzyas:

Okay! Thanks @Convert

Convert:

Happy to help! :) and sorry about before I read 3.50 while it was 3.30

Aricuzyas:

:)